# Consistency in Language Models: Current Landscape, Challenges, and Future Directions

### v1

@@ Abstract
The hallmark of effective language use lies in consistency—expressing similar meanings in similar contexts and avoiding contradictions. While human communication naturally demonstrates this principle, state-of-the-art language models struggle to maintain reliable consistency across different scenarios. This paper examines the landscape of consistency research in AI language systems, exploring both formal consistency (including logical rule adherence) and informal consistency (such as moral and factual coherence). We analyze current approaches to measure aspects of consistency, identify critical research gaps in standardization of definitions, multilingual assessment, and methods to improve consistency. Our findings point to an urgent need for robust benchmarks to measure and interdisciplinary approaches to ensure consistency in the application of language models on domain-specific tasks while preserving the utility and adaptability.
@@ 1 Introduction
Consistency—broadly defined as using language similarly in similar settings or avoiding contradictions when using language—is among the most important forms of generalization in the use of language. This ability to maintain consistent outputs is essential for building reliable AI systems that users can trust and depend on. Consistency serves as both a natural expectation users have when interacting with language technologies and a prerequisite for deploying these models in high-stakes domains (Elazar et al., 2021; Jang et al., 2022; Kim et al., 2025). However, the most advanced large language models (LLMs) struggle with this and demonstrate multiple examples of inconsistent behavior (Elazar et al., 2021; Raj et al., 2025). Although such examples have been documented in multiple studies, there are no standard approaches to assessing model consistency. As such, there is an ongoing risk of overestimating the performance of state-of-the-art models, as well as of underestimating the risks and potential harms elicited by such models.
@@
There have been some early attempts to analyze the behavior of models in terms of consistency, to understand the roots of inconsistency, and to discover ways to mitigate inconsistency and/or to use inconsistent outputs to improve the model. However, research on language model consistency faces multiple challenges, starting from a lack of agreement on terminology and evaluation methods, followed by limitations related to data and models. In this paper, we present a review of current research on consistency in language models, highlight the most pressing challenges, and provide recommendations for future research directions. For greater specificity, we restrict our attention to text-only language models, on which an overwhelming majority of existing research on consistency has been based. See Appendix A for references to some research on multimodal consistency.
2 A Review of Consistency Research in ML
@@
Consistency in language models connects to several critical areas in AI research: hallucinations (enerating made-up information that may contradict known facts, factuality (agreement with real-world knowledge), misinformation (false claims misleading users), and reasoning (logical coherence across statements). As such, consistency evaluation forms a crucial component of foundation model assessment frameworks. This review surveys literature on consistency in language models from 2019 to 2025, focusing on peer-reviewed publications from major AI conferences and influential preprints that explicitly address consistency through evaluation, enhancement techniques, or theoretical analysis.
2.1 Terminology and Consistency Types
@@
The terminology used to describe the consistency of language models is often confusing, as there is not a single, commonly agreed-upon definition of consistency. Authors either come up with their own definition of the concept that aligns best with the specifics of the work they focus on, or use an overly broad definition, or sometimes just omit defining the term altogether. As a result, existing research studies present multiple narrowly focused definitions of consistency that often cover very different aspects of model behavior and sometimes even contradict each other.
@@
Despite disagreements about consistency definitions in previous studies of language models, all of them, in one way or another, relate to the behavior of these models. Given this interest in the behavior of models and the implications for potential model applications, in this paper we limit the otherwise broader concept of consistency to behavioral consistency. Behavioral consistency has been thoroughly analyzed in the field of human personality psychology, where it refers to adhering to the same principles in behavioral patterns across a large variety of situations and sometimes over many years (Funder & Colvin, 1991). In psychology, behavioral consistency is tightly linked with the predictability of behavior, which is equally important for the applications of language models. Based on how behavioral consistency is approached in the literature, in this paper, we further categorize the different types of consistency into two large groups: logical/formal and non-logical/informal.
2.1.1 Logical / formal consistency
@@
Logical consistency in relation to large language models was introduced by Jang et al. (2022), as an ability of the model to make decisions without logical contradiction. The rules and principles of formal logic are applied to assess the behavior of a model in a methodical way, making the evaluation standardized and easy to follow. Based on these principles, Jang et al. (2022) classify consistency into negational, symmetric, transitive, and additive types. Negational consistency follows the logical negation property ($p$ is true $\Leftrightarrow\neg{p}$ is false), i.e. LM’s predictions should be opposite for texts with the opposite meanings. Symmetric consistency follows the rule $f(x,y)=f(y,x)$ and implies that LM ’s predictions should be invariant to the input text swap. Transitive consistency can measure deductive reasoning ability and follows the property of transitive inference, represented as $X\rightarrow Y\land Y\rightarrow Z$ then $X\rightarrow Z$. This type of consistency was analyzed in NLI tasks Li et al. (2019) as well as in question answering Asai & Hajishirzi (2020b); Mitchell et al. (2022).
@@
Semantic consistency, another subpart of the Jang et al. (2022) definition, is one of the most frequently used concepts in existing consistency research studies. The idea of semantic consistency is derived from the semantic equivalence property, represented as $f(X)=f(Y)$ if $X$ and $Y$ mean the same. It was explored by Elazar et al. (2021) as the ability of a model to make consistent decisions in semantically equivalent contexts, and a very similar understanding of the concept was later used in multiple studies Raj et al. (2023; 2022); Ohmer et al. (2024).
2.1.2 Nonlogical / informal consistency.
@@
The nonlogical, or informal, type of consistency covers all the other definitions that do not follow the rules of formal logic. For example, Bonagiri et al. (2024) highlight the importance of moral consistency, as the ability to preserve non-contradictory moral values across different types of situations (Arvanitis and Kalliris, 2020; Marcus, 1980), as a part of LLM alignment. Their approach consists of generating semantically equivalent scenarios and employing consistency checks to see if a target LLM gets the same Semantic Graph Entropy (SaGE) score while responding to these scenarios. To assess the reliability of SaGE, they use human evaluation. Jain et al. 2024 investigate norm inconsistency in LLMs, which they define as condition where LLMs apply different norms in similar situations. They specifically focus on applications of models in high-risk domains.
@@
Informational and/or factual consistency is another subpart of the Jang et al. (2022) definition and is a term frequently used in consistency research work. Manakul et al. (2023) use the term informational consistency, without explaining or defining it further, in developing a method of fact-checking the responses of black-box models. The term factual consistency is often used in the context of automatic summarization Wang et al. (2020). Factual inconsistency is often referred to as hallucinations and/or faithfulness, i.e. models generating new information that contradicts the source document (e.g. Tam et al, 2023, Maynez et al., 2020). Definitions of factual consistency are often not clearly specified and instead replaced with human annotations, as in e.g. Tam et al. (2023). [^1] There is a case to be made that factual and moral consistency are better construed as a logical property. For instance, factual consistency could be construed as the absence of contradictions between a model’s outputs, and factually true statements.
@@
Some more recent work clarifies the concept of consistency. A recent study by Parcalabescu & Frank (2024) on natural language explanations contrasts faithfulness and self-consistency. Self-consistency examines whether similar inputs produce consistent explanations—essentially measuring explanation stability across input variations. Faithfulness, meanwhile, evaluates whether the explanation behind a certain model-generated answer accurately reflects the model’s reasoning process to come up with that answer. While related, they involve different evaluation approaches. Self-consistency requires testing multiple input variations (which may not generalize well across datasets) and does not necessarily involve checking for accuracy of the explanations. On the other hand, faithfulness focuses on the accuracy of individual explanations without such constraints.
2.2 Analyzed Downstream Tasks
@@
Of these, a slim majority of studies on language model consistency investigate well-established NLP tasks. The tasks most commonly analyzed are question answering (QA)Mündler et al. (2024); Raj et al. (2023; 2022); Berglund et al. (2024); Li et al. (2023); Wang et al. (2020); Asai & Hajishirzi (2020a); Liu et al. (2024), summarization West et al. (2024); Cui et al. (2024); Tam et al. (2023); Wang et al. (2020), natural language inference (NLI) Jang et al. (2022); Jang & Lukasiewicz (2023); Camburu et al. (2020); West et al. (2024); Dziri et al. (2019) and reasoning Zhang et al. (2024b); Liu et al. (2024); Chen et al. (2024); Wang et al. (2023) Roughly a third of existing studies do not rely on standard NLP tasks. These studies typically use custom tasks, such as generating continuations of sentences from Wikipedia Mündler et al. (2024). A small number of studies use an approach tailored to specific use cases, e.g. measuring the accuracy of stock price predictions based on textual information such as earnings calls and news articles Yang et al. (2023).
2.3 Size and Availability of Testing Data
@@
The number of testing samples varies quite substantially across different studies, from a few hundred to tens of thousands of samples. One standard approach to creating a test dataset for consistency evaluation is to use one or several of the existing standard benchmarks and modify them according to a set of rules or templates, so that consistency assessment is possible (Jang et al., 2022; Fierro & Søgaard, 2022). Another approach is to enhance existing benchmarks with either human- or LLM-generated annotations (Liu et al., 2023). When evaluating semantic consistency, a common way is to create paraphrases of an existing dataset using automatic parahrasing methods (Bonagiri et al., 2024) and/or human annotators (Elazar et al., 2021). The majority of resulting testing datasets are shared publicly for replicability of experiments, although in some cases the authors rely on the description of the dataset creation process and do not provide access to the actual dataset.
2.4 Evaluated Models
@@
More than two-thirds of the studies on language model consistency in our literature review use generative, transformer-based LMs with decoder-only or encoder-decoder architectures, such as the GPT and OPT series models, BART, and T5 Jang et al. (2022); Li et al. (2023); Jang & Lukasiewicz (2023); West et al. (2024); Berglund et al. (2024); Raj et al. (2023; 2022); Manakul et al. (2023); Mündler et al. (2024); Zhang et al. (2024b); Liu et al. (2024); Zhang et al. (2024a); Cheng et al. (2024); Cui et al. (2024); Tam et al. (2023); Wang et al. (2023); Chen et al. (2024); Wang et al. (2020); Nie et al. (2021); Qin et al. (2021). Model sizes for these studies range from a few billion parameters to hundreds of billions of parameters or more. Many of these studies also consider other types of models. Slightly more than half of the papers we reviewed test proprietary models such as GPT-4 (Li et al., 2023; Jang & Lukasiewicz, 2023; West et al., 2024; Berglund et al., 2024; Mündler et al., 2024; Zhang et al., 2024b; Liu et al., 2024; Zhang et al., 2024a; Cui et al., 2024; Chen et al., 2024; Cheng et al., 2024), whose exact sizes have not been publicly disclosed, but in some cases are rumored to exceed a trillion parameters. About a quarter of the studies Jang et al. (2022); Elazar et al. (2021); Asai & Hajishirzi (2020a); Yang et al. (2023); Nie et al. (2021); Qin et al. (2021) focus on encoder-only, BERT-style models such as BERT, RoBERTa, and ALBERT.
2.5 Evaluation Metrics
@@
Metrics to measure different notions of consistency typically depend on pairwise similarity metrics. They compute intrinsic base metrics such as BERTScore, ROUGE, Entailment, or Contradiction for pairs of outputs given similar inputs and/or context, and aggregate over multiple such pairs. In earlier studies, the base metrics were based on token-matching similarities (Elazar et al., 2021). Later papers graduated to notions of semantic similarity that are robust to syntactic variations that can change the wording or structure of a phrase of text while keeping the meaning the same or similar (Raj et al., 2022; 2023; Rabinovich et al., 2023; Manakul et al., 2023). Aggregation of a metric across a set of pairs is typically done using simple averaging, with the exception of Mündler et al. (2024), which uses sequential aggregation of contradiction scores to detect factual consistency, and (Raj et al., 2023; Kuhn et al., 2023) who use semantic entropy across the full set of outputs.
2.6 Other Important Aspects
@@
In spite of recent attention, certain important aspects of consistency remain underresearched. We specifically call out two such directions. First, recent consistency research tends to focus excessively on consistency in generations at decoding time. In this process it ignores encoder-only models (e.g. those based on BERT and RoBERTa), and how (in)consistent inputs shape the performance on downstream standard NLP tasks such as sentiment prediction. Another direction currently underexplored is adversarial attacks to degrade consistency. While there is extensive research on adversarial robustness (e.g. the AdvGLUE benchmark(Wang et al., 2022)) and jailbreaks, there are very few studies exploring how inconspicuous or subtle manipulation of prompts can lead to inconsistent LLM responses (Lin et al., 2024a), and we do not yet fully understand how much malicious perturbations coupled with slightly different input text can degrade the quality of the downstream output.
@@
The availability of model weights and training datasets—which allow for stronger transparency and reproducibility—aid in investigating the root causes of inconsistency in LLMs. As shown by Lin et al. (2024b), analyzing the internal state of the model can improve the transparency of the model and lay the foundation for mitigating model hallucinations and inconsistencies. Not only closed-weight models, but also unpublished source code and unpublished datasets make it nearly impossible to reproduce some existing publications, making it very hard to verify the claims and findings stated in the publications (Semmelrock et al., 2023).
3 Discussion and Recommendations for Future Research
3.1 Standardization of Consistency Research
In section 2, we reviewed existing consistency research and showed that some types of consistency, such as e.g. transitive or symmetric, and non-English languages are underrepresented in current literature, while other aspects, such as semantic consistency of English-only text-to-text models, are relatively well researched. We also showed that there is very little consensus on the general definition of consistency and the use of evaluation metrics. Due to the lack of standard evaluation processes and benchmarks, the majority of existing work has been performed on individually created datasets that are not always openly shared. This lack of standard evaluation methods makes it difficult to compare the results of individual consistency studies, replicate previous works, and as a result, understand the progress of model development when it comes to consistency evaluation.
3.2 Multilingual Consistency

### v2

@@ Abstract
The hallmark of effective language use lies in consistency: expressing similar meanings in similar contexts and avoiding contradictions. While human communication naturally demonstrates this principle, state-of-the-art language models (LMs) struggle to maintain reliable consistency across task- and domain-specific applications. Here we examine the landscape of consistency research in LMs, analyze current approaches to measure aspects of consistency, and identify critical research gaps. Our findings point to an urgent need for quality benchmarks to measure and interdisciplinary approaches to ensure consistency while preserving utility.
@@ 1 Introduction
Consistency—broadly defined as using language similarly in similar settings or avoiding contradictions when using language—is among the most important forms of generalization in the use of language. This ability to maintain consistent outputs is essential for building reliable AI systems that users can trust and depend on. Consistency is both a natural expectation that users have when interacting with language technologies and a prerequisite to deploying them in high-stakes domains (Elazar et al., 2021; Jang et al., 2022; Kim et al., 2025). However, most advanced large language models (LLMs) struggle with consistency and frequently demonstrate inconsistent behavior (Elazar et al., 2021; Raj et al., 2025). Although such examples have been documented in multiple studies, there are no standard approaches to assessing model consistency. As such, there is an ongoing risk of overestimating the performance of state-of-the-art models, as well as of underestimating the risks and potential harms elicited by them.
@@
Despite early attempts to measure and enhance the consistency of language models (LM) and to understand the roots of inconsistency, research on this topic faces multiple challenges. These include a lack of agreement on terminology and evaluation metrics, and limitations on data and model availability. In this paper, we present a review of current research on consistency in LMs, highlight the most pressing challenges, and provide recommendations for future research. We restrict our attention to text-only LMs, which a majority of existing research is based on. See Appendix A for a brief discussion on multimodal consistency.
2 A Review of Consistency Research
@@
Consistency has connections to critical areas in AI research: hallucination (generating made-up information contradicting references), factuality (agreement with real-world knowledge), misinformation (false claims misleading users), and reasoning (logical coherence across statements). We survey literature on consistency in LMs from 2019 to 2025, focusing on peer-reviewed publications and influential preprints that explicitly address consistency metrics, theory, and enhancement.
Terminology
@@
The terminology used to describe the consistency of LMs is often confusing, as there is not a single, commonly agreed-upon definition of consistency. Authors either come up with their own definition of the concept that aligns best with the specifics of the work they focus on, or use an overly broad definition, or sometimes just omit defining the term altogether. As a result, existing studies present multiple narrowly focused definitions of consistency that often cover very different aspects of model behavior and sometimes even contradict each other.
@@
Given this interest in model behavior and the implications for potential model applications, in this paper, we limit the otherwise broader concept of consistency to behavioral consistency. In psychology, behavioral consistency is closely related to the predictability of behavior, which is equally important for the applications of LMs. Based on how behavioral consistency is approached in the literature, we categorize the different types of consistency into two large groups: logical/formal and nonlogical/informal.
@@
Logical consistency in LLMs was introduced by Jang et al. (2022), as the ability of the model to make decisions without logical contradiction. The rules and principles of formal logic are applied to assess the behavior of a model in a methodical way, allowing for standardized and intuitive measurement. Based on these principles, Jang et al. (2022) classified consistency into negational, symmetric, transitive, and additive types. Negational consistency follows the logical negation property ($p$ is true $\Leftrightarrow\neg{p}$ is false), i.e. LM’s predictions should be opposite for texts with the opposite meanings. Symmetric consistency follows the rule $f(x,y)=f(y,x)$ and implies that the predictions of an LM should be invariant to the input text swap. Transitive consistency can measure deductive reasoning ability and follows the property of transitive inference, represented as $X\rightarrow Y\land Y\rightarrow Z$ then $X\rightarrow Z$. This type of consistency was analyzed in natural language inference (NLI) tasks Li et al. (2019) and question-answering (Q&A) Asai & Hajishirzi (2020b); Mitchell et al. (2022).
@@
Semantic consistency, another subpart of the Jang et al. (2022) definition, is one of the most widely used concepts in existing consistency research studies. The idea of semantic consistency is derived from the semantic equivalence property, represented as $f(X)=f(Y)$ if $X$ and $Y$ mean the same. Elazar et al. (2021), and multiple studies later on Raj et al. (2022); Ohmer et al. (2024), explored this as the ability of a model to make consistent decisions in semantically equivalent contexts.
@@
Nonlogical or informal consistency covers all the other definitions that do not follow the rules of formal logic. For example, Bonagiri et al. (2024) highlight the importance of moral consistency, as the ability to preserve noncontradictory moral values across different situations Arvanitis & Kalliris (2020); Marcus (1980), in LLM alignment. Their approach consists of generating semantically equivalent scenarios and employing consistency checks to see if a target LLM gets the same Semantic Graph Entropy (SaGE) score while responding to these scenarios. Jain et al. (2025) investigated norm inconsistency, defined as the condition in which LLMs apply different norms in similar situations, on applying LLMs in high-risk domains.
@@
Informational and/or factual consistency is another subpart of the Jang et al. (2022) definition frequently used in consistency research. Manakul et al. (2023) used the term informational consistency, without explaining or defining it further, to develop a method for fact-checking the responses of black-box models. The term factual consistency is often used in the context of automatic summarization Wang et al. (2020). Factual inconsistency is often referred to as hallucinations and/or faithfulness, i.e., models that generate new information that contradicts the source document Tam et al. (2023); Maynez et al. (2020). Definitions of factual consistency are often not clearly specified, and instead are replaced with human annotations.
@@
A recent study Parcalabescu & Frank (2024) on natural language explanations contrasts faithfulness and self-consistency. Self-consistency examines whether similar inputs produce consistent explanations—essentially measuring explanation stability across input variations. Faithfulness, meanwhile, evaluates whether the explanation behind a certain model-generated answer accurately reflects the model’s reasoning process to come up with that answer. While related, they involve different evaluation approaches. Self-consistency requires testing multiple input variations (which may not generalize well across datasets) and does not necessarily involve checking for accuracy of the explanations. On the other hand, faithfulness focuses on the accuracy of individual explanations without such constraints.
Analyzed Tasks
@@
A slim majority of studies on LM consistency investigate well-established NLP tasks. Most commonly analyzed tasks include Q&A Mündler et al. (2024); Raj et al. (2022); Berglund et al. (2024); Li et al. (2023); Wang et al. (2020); Asai & Hajishirzi (2020a), summarization West et al. (2024); Cui et al. (2024); Tam et al. (2023); Wang et al. (2020), NLI Jang et al. (2022); Jang & Lukasiewicz (2023); Camburu et al. (2020); West et al. (2024); Dziri et al. (2019) and reasoning Zhang et al. (2024b); Liu et al. (2024b); Chen et al. (2024); Wang et al. (2023) Approximately a third of existing studies do not rely on standard NLP tasks, usually using custom tasks such as generating continuations of sentences from Wikipedia Mündler et al. (2024). A small number of studies employ use-case specific approaches, for example, measuring stock price prediction accuracy based on textual information such as earnings calls and news articles Yang et al. (2023).
Dataset Size and Availability
@@
The number of testing samples varies substantially across different studies, from a few hundred to tens of thousands. One standard approach to creating a test dataset for measuring consistency is to multiply the prompts in one or more existing benchmarks using perturbation rules or prompt templates (Jang et al., 2022; Fierro & Søgaard, 2022). Another approach is to enhance existing benchmarks with human- or LLM-generated annotations (Liu et al., 2023). To do this, a common method is to create paraphrases of an existing dataset using automatic parahrasing methods (Bonagiri et al., 2024) and/or human annotators (Elazar et al., 2021). The majority of testing datasets are shared publicly, although in some cases the authors only describe the dataset creation process without providing access to the actual dataset.
Evaluated Models
@@
More than two-thirds of the studies we examined use transformer-based generative LMs with decoder-only or encoder-decoder architectures, such as the GPT and OPT series models, BART, and T5 Jang et al. (2022); Li et al. (2023); Jang & Lukasiewicz (2023); West et al. (2024); Berglund et al. (2024); Raj et al. (2022); Manakul et al. (2023); Mündler et al. (2024); Zhang et al. (2024b); Liu et al. (2024b); Zhang et al. (2024a); Cheng et al. (2024); Cui et al. (2024); Tam et al. (2023); Wang et al. (2023); Chen et al. (2024); Wang et al. (2020); Nie et al. (2021). The parameter sizes for the models tested range from a few billion to hundreds of billion. Slightly more than half of the papers test proprietary models such as GPT-4 (Li et al., 2023; Jang & Lukasiewicz, 2023; West et al., 2024; Berglund et al., 2024; Mündler et al., 2024; Zhang et al., 2024b; Liu et al., 2024b; Zhang et al., 2024a; Cui et al., 2024; Chen et al., 2024), whose exact sizes have not been publicly disclosed but in some cases are rumored to exceed a trillion parameters. Some studies also consider other types of LMs: about a quarter of papers Jang et al. (2022); Elazar et al. (2021); Asai & Hajishirzi (2020a); Yang et al. (2023); Nie et al. (2021); Qin et al. (2021) focus on encoder-only, BERT-style models such as BERT, RoBERTa, and ALBERT.
Evaluation of Consistency
Consistency evaluation typically uses two approaches: (1) input-based sampling, creating paraphrases or equivalent prompts to test consistent responses to similar inputs, or (2) output-based sampling, generating multiple outputs from identical inputs. Output-based sampling with high temperature may artificially inflate inconsistency by forcing models to sample normally-avoided tokens, potentially misrepresenting model behavior.
@@
Metrics to measure different notions of consistency typically depend on pairwise similarity metrics. They compute base metrics such as BERTScore, ROUGE, Entailment, or Contradiction for pairs of outputs given similar inputs and/or context, and aggregate over multiple pairs. In earlier studies, the base metrics were based on token-matching similarities (Elazar et al., 2021). Later papers graduated to notions of semantic similarity that are robust to syntactic variations that can change the wording or structure of a phrase of text while keeping the meaning the same or similar (Raj et al., 2022; Rabinovich et al., 2023; Manakul et al., 2023). Aggregation of a metric across pairs is typically done by simple averaging, with the exception of Mündler et al. (2024), which uses sequential aggregation of contradiction scores to measure factual consistency, and Raj et al. (2025); Kuhn et al. (2023) who use semantic entropy across the entire set of outputs.
Challenges
@@
Two important aspects of consistency remain underresearched. First, current work tends to focus excessively on consistency in generations at decoding time. In this process, it ignores encoder-only models and how (in)consistent inputs shape the performance on downstream standard NLP tasks like sentiment prediction. Another underexplored direction is adversarial attacks to degrade consistency. Despite extensive research on adversarial robustness (e.g. the AdvGLUE benchmark (Wang et al., 2022)) and jailbreaks, very few studies explore how inconspicuous or subtle manipulation of prompts can lead to inconsistent LLM responses (Lin et al., 2024a). We do not yet fully understand how much malicious perturbations coupled with slightly different input text can degrade output quality.
@@
The availability of model weights and training datasets—allowing stronger transparency and reproducibility—aid in investigating the root causes of inconsistency. Lin et al. (2024b) showed that analyzing the internal state of the model can improve the transparency of the model and lay the foundation for mitigating hallucinations and inconsistencies. Not only closed-weight models, but also unpublished source code and datasets make it nearly impossible to reproduce and verify claims and findings of some existing publications (Semmelrock et al., 2023).
3 Discussion and Recommendations
As mentioned earlier, we need standardization of terms and definitions for a better understanding of the progress of consistent language model development. Beyond this, we recommend the following focus areas for future research.
Multilingual Consistency